OpenStudy (anonymous):

List all possible rational zeros of the function: 4x^3 -3x^2 +4x -8

OpenStudy (anonymous):

make all possible fractions where numerator divides 8 and denominator divides 4 (constant and leading coefficient respectively)

OpenStudy (anonymous):

divisors of 8 are \[\pm1,\pm2,\pm4,\pm8\]

OpenStudy (anonymous):

then i guess you need to add some more like \[\pm\frac{1}{2}\]and \[\pm\frac{1}{4}\] i think that is all of them

OpenStudy (anonymous):

then i guess you need to add some more like \[\pm\frac{1}{2}\]and \[\pm\frac{1}{4}\] i think that is all of them

OpenStudy (anonymous):

Thank you!